# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 3.3.1    First excitation shell
We start with the first excitation shell $\text{N}=1$, which appears in the negative-parity ($P=-1$) sector shown in the left panel in Fig. 3.3. There is a low-mass cluster of two nearly mass degenerate $N^{*}$’s around 1.5 GeV, the $N(1535)\nicefrac{{1}}{{2}}^{-}$ and $N(1520)\nicefrac{{3}}{{2}}^{-}$, where the former is the nucleon’s parity partner. In the $SU(6)$ assignment these correspond to $(\mathbf{D},L^{P},S)=(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$, i.e., their wave functions have $SU(6)$ dimensionality $\mathbf{70}$ and thus mixed symmetry, and they carry total quark spin $S=\nicefrac{{1}}{{2}}$ and orbital angular momentum $L=1$ which are coupled to $J=\nicefrac{{1}}{{2}}$ and $\nicefrac{{3}}{{2}}$. The next cluster around 1.65 GeV contains a $N^{*}$ triplet, the $N(1650)\nicefrac{{1}}{{2}}^{-}$, $N(1700)\nicefrac{{3}}{{2}}^{-}$, and $N(1675)\nicefrac{{5}}{{2}}^{-}$ forming the $N(\mathbf{70},1^{-},\nicefrac{{3}}{{2}})$, and a $\Delta^{*}$ pair, the $\Delta(1620)\nicefrac{{1}}{{2}}^{-}$ and $\Delta(1700)\nicefrac{{3}}{{2}}^{-}$ forming the $\Delta(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$. Note that the first shell does not permit a $\Delta$ triplet.
The c.o.g. masses of these states are collected in Fig. 3.5. The small mass splittings within the two spin doublets and the spin triplet imply that the $\bm{L}\cdot\bm{S}$ coupling is small. The mass-square difference between the $N(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$ and the mean value of the five remaining states is 0.56(7) GeV2. In principle, the states with identical quantum numbers can mix, but the mixing angle is small: Experimentally and in calculations, the mixing angle is determined to be $-31,7^{\circ}$ for the two $\nicefrac{{1}}{{2}}^{-}$ and +10 for the $\nicefrac{{3}}{{2}}^{-}$ states [534]. Here, we caution the reader: the mixing angle was derived assuming a small branching ratios of about 1% for $N(1650)\nicefrac{{1}}{{2}}^{-}\to N\eta$ from unreliable data on $\pi^{-}p\to\eta p$; the later determination was reported in Ref. [394]. Recent determinations using precise data on $\eta$ photoproduction find, however, a branching ratio of 28(11)[131] and 33(4)[93].
@@
Figure 3.5: Squared masses of stretched $N^{*}$’s (a) and $\Delta^{*}$’s (b) as a function of their spin $J$. c: C.o.g. masses (in MeV) for the baryon excitations in Fig. 3.2 up to the second excitation shell. d: Mass splittings in the second and third excitation shell expected in the quark model, see Refs. [534, 544, 545].
@@ 3.3.2    Second excitation shell
$(\mathbf{56},0^{+})$: The so-called radial excitations of the nucleon and $\Delta(1232)$ belong to the 56-plet, so they have symmetric orbital and spin-flavor wave functions and carry orbital angular momentum $L=0$. The low mass of 1366(3) MeV for the Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ compared to the $N(1535)\nicefrac{{1}}{{2}}^{-}$ has led to speculations concerning its nature [546]. Its squared mass is 0.98(1) GeV2 above $M^{2}_{N}$. The $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ with a pole mass of 1550(15) MeV is the analogue of the Roper and was nicknamed ‘Doper’ by Nefkens (who also introduced the names ‘Loper’ for $\Lambda(1600)\nicefrac{{1}}{{2}}^{+}$, ‘Soper’ for $\Sigma(1660)\nicefrac{{1}}{{2}}^{+}$ and ‘Xoper’ for the not yet identified spin-$\nicefrac{{1}}{{2}}^{+}$ $\Xi$ resonance in his talk at MENU2004 [547]). The $\Delta(1600)\nicefrac{{3}}{{2}}^{+}-N(1440)\nicefrac{{1}}{{2}}^{+}$ mass-square difference of 0.54(4) GeV2 is compatible with the $\Delta(1232)-N(940)$ mass-square difference of 0.58 GeV2.
@@
$(\mathbf{70},0^{+})$: The remaining $\text{N}=2$ mass region is rather densely packed with baryon excitations. The $N(1710)\nicefrac{{1}}{{2}}^{+}$ has a pole mass of 1696(10) MeV. Given that its squared mass splitting with the Roper resonance of 1.01(4) GeV2 is of comparable size as that of the Roper versus the nucleon, it could be interpreted as the second radial excitation of the nucleon in a 56-plet. In quark models, however, the $N(1710)\nicefrac{{1}}{{2}}^{+}$ is found in a 70-plet, which has a spatial wave functions with mixed symmetry. In that case, both $\rho$ and $\lambda$ oscillators carry one angular momentum excitation and we expect a sizable fraction of all decays to proceed via an intermediate excitation, i.e., via $N(1440)\nicefrac{{1}}{{2}}^{+}$ or $N\sigma$. In the RPP these two decay modes are not listed as seen, but in a recent analysis the former branching ratio is reported with a frequency 22(6)and the latter with 11(4)[474, 475]. Thus, in agreement with the quark model interpretation we consider the $N(1710)\nicefrac{{1}}{{2}}^{+}$ as the $N(\mathbf{70},0^{+},\nicefrac{{1}}{{2}})$ state. We then further expect a nucleon spin partner $N(\mathbf{70},0^{+},\nicefrac{{3}}{{2}})$ with $J=\nicefrac{{3}}{{2}}^{+}$ in this multiplet. This could be the $N^{\prime}(1720)\nicefrac{{3}}{{2}}^{+}$ suggested in a common photo- and electroproduction analysis of $p\pi^{+}\pi^{-}$ [310], which is not included in the RPP 2024. The remaining decuplet state $\Delta(\mathbf{70},0^{+},\nicefrac{{1}}{{2}})$ could be the 1* resonance $\Delta(1750)\nicefrac{{1}}{{2}}^{+}$ with a mass of 1770(30) MeV.
@@
$(\mathbf{56},2^{+})$: In the same mass region there are several further states: a spin doublet $N(\mathbf{56},2^{+},\nicefrac{{1}}{{2}})$ containing the $N(1720)\nicefrac{{3}}{{2}}^{+}$ and $N(1680)\nicefrac{{5}}{{2}}^{+}$ with c.o.g. mass 1689(13) MeV, and a spin quartet $\Delta(\mathbf{56},2^{+},\nicefrac{{3}}{{2}})$ with the $\Delta(1910)\nicefrac{{1}}{{2}}^{+}$, $\Delta(1920)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1905)\nicefrac{{5}}{{2}}^{+}$ and $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$ and c.o.g. mass 1859(20) MeV. We note that the $\Delta(1905)\nicefrac{{5}}{{2}}^{+}$ has an unexpectedly low pole mass of 1795(25) MeV. The mass square from $\Delta(\mathbf{56},2^{+},\nicefrac{{3}}{{2}})$ to $N(\mathbf{56},2^{+},\nicefrac{{1}}{{2}})$ is 0.60(8) GeV2 and 2×2$\times$2 × 1.00(8) GeV2 to the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$.
@@ 3.3.3    The third shell
$(\mathbf{56},1^{-})$: There are five negative-parity states with masses that are difficult to reconcile in quark models. In [548, 549], the lowest-mass $N^{*}$ doublet in the third shell is expected at about 1950 MeV and a $\Delta^{*}$ triplet at around 2100 MeV. Experimentally, we find a cluster of $N^{*}$’s forming a spin-doublet, the $N(1895)\nicefrac{{1}}{{2}}^{-}$ and $N(1875)\nicefrac{{3}}{{2}}^{-}$ with c.o.g. mass 1868(23) MeV, and the spin triplet $\Delta(1900)\nicefrac{{1}}{{2}}^{-}$, $\Delta(1940)\nicefrac{{3}}{{2}}^{-}$, and $\Delta(1930)\nicefrac{{5}}{{2}}^{-}$ with c.o.g. mass 1858(63) MeV. This would fit with the $N(\mathbf{56},1^{-},\nicefrac{{1}}{{2}})$ and $\Delta(\mathbf{56},1^{-},\nicefrac{{3}}{{2}})$ assignment. To explain its low mass, one of these states, the $N(1895)\nicefrac{{1}}{{2}}^{-}$, has been interpreted as a dynamically generated state with a two-pole structure [550, 551].
@@ 3.3.4    Beyond the third shell
In the negative-parity sector, the $\Delta(2350)\nicefrac{{5}}{{2}}^{-}$ and $\Delta(2400)\nicefrac{{9}}{{2}}^{-}$ could be naturally assigned to the $\Delta(\mathbf{56},3^{-},\nicefrac{{3}}{{2}})$ multiplet in the third excitation shell (see Fig. 3.2). However, the former 1* resonance is observed at 2400(125) MeV [359], while for the latter 2* resonance two measurements have been reported, 2260(60) MeV [359] and 2458(2) MeV [458]. The unweighted mean of these three values is 2375(100) MeV, which is considerably higher than the masses of the other states in the third shell. Moreover, the masses of these two resonances fall onto a Regge trajectory with an offset compared to main trajectory given by the $\Delta(1600)\nicefrac{{3}}{{2}}^{+}-\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ squared mass separation, as shown in the right panel of Fig. 3.6. We thus prefer an interpretation of these states as members of $\Delta(\mathbf{56},3^{-},\nicefrac{{3}}{{2}})$ in the fifth excitation shell ($\text{N}=5$). For the same reasons one can assign the 3* resonance $N(2600)\nicefrac{{11}}{{2}}^{-}$ to the multiplet $N(\mathbf{70},5^{-},\nicefrac{{1}}{{2}})$ in the fifth shell and the $\Delta(2750)\nicefrac{{13}}{{2}}^{-}$ with mass 2794(80) MeV to $\Delta(\mathbf{56},5^{-},\nicefrac{{3}}{{2}})$ in the seventh shell.
Concerning the positive-parity resonances, the $\Delta(2420)\nicefrac{{11}}{{2}}^{+}$, $\Delta(2300)\nicefrac{{9}}{{2}}^{+}$ and $\Delta(2390)\nicefrac{{7}}{{2}}^{+}$ have a c.o.g. mass of 2360(100) MeV and can be combined into a multiplet $\Delta(\mathbf{56},4^{+},\nicefrac{{3}}{{2}})$ in the fourth shell, where the $J=\nicefrac{{5}}{{2}}^{+}$ state is missing. In the nucleon sector, the RPP lists the 4* resonance $N(2220)\nicefrac{{9}}{{2}}^{+}$; the $N(2200)\nicefrac{{7}}{{2}}^{+}$ suggested by Hunt and Manley [394] could serve as its spin partner to form a $N(\mathbf{56},4^{+},\nicefrac{{1}}{{2}})$ in the fourth shell. Finally, the $N(2700)\nicefrac{{13}}{{2}}^{+}$ and $\Delta(2950)\nicefrac{{15}}{{2}}^{+}$ are the highest-mass states included in the RPP; the Regge trajectories in Fig. 3.6 suggest to place the former in the sixth shell with $N(\mathbf{56},6^{+},\nicefrac{{1}}{{2}})$ and the latter in the eight shell with $\Delta(\mathbf{56},6^{+},\nicefrac{{3}}{{2}})$.
Table 3.2: Number of expected and (in parentheses) reported $N^{*}$ and $\Delta^{*}$ resonances with $J^{P}=\nicefrac{{1}}{{2}}^{\pm},\cdots,\nicefrac{{9}}{{2}}^{\pm}$ in the first three excitation shells. The $N^{\prime}(1720)\nicefrac{{3}}{{2}}^{+}$ suggested in electroproduction  [310] and $N(1965)\nicefrac{{3}}{{2}}^{+}$ suggested in [474, 475] are not included in the counting.
$N$ Shell $P$ $J=\frac{1}{2}$ $J=\frac{3}{2}$ $J=\frac{5}{2}$ $J=\frac{7}{2}$ $J=\frac{9}{2}$
1st -- 2 (2) 2 (2) 1 (1)
@@
3rd -- 3 (2) 5 (1) 4 (2) 2 (1) 1 (0)
@@
$X$ $Y$ $\delta M^{2}$ [GeV2]
$N$ $N(1440)\frac{1}{2}^{+}$ $0.98\pm 0.01$
$N(1440)\frac{1}{2}^{+}$ $N(1710)\frac{1}{2}^{+}$ 1.01$\pm$0.04
$\Delta(1232)\frac{3}{2}^{+}$ $\Delta(1600)\frac{3}{2}^{+}$ 0.94$\pm$0.05
$\Lambda$ $\Lambda(1600)\frac{1}{2}^{+}$ 1.16$\pm$0.03
$\Sigma$ $\Sigma(1660)\frac{1}{2}^{+}$ 1.26$\pm$0.06
$\Delta(1620)\frac{1}{2}^{-}$ $\Delta(1900)\frac{1}{2}^{-}$ 0.74$\pm$0.08
$\Delta(1700)\frac{3}{2}^{-}$ $\Delta(1940)\frac{3}{2}^{-}$ 1.36$\pm$0.17

### v4

@@ 3.3.1   First excitation shell
We start with the first excitation shell $\text{N}=1$, which appears in the negative-parity (P=1P=-1) sector shown in the left panel in Fig. 3.3. There is a low-mass cluster of two nearly mass degenerate $N^{*}$’s around 1.5 GeV, the $N(1535)\nicefrac{{1}}{{2}}^{-}$ and $N(1520)\nicefrac{{3}}{{2}}^{-}$, where the former is the nucleon’s parity partner. In the $SU(6)$ assignment these correspond to $(\mathbf{D},L^{P},S)=(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$, i.e., their wave functions have $SU(6)$ dimensionality $\mathbf{70}$ and thus mixed symmetry, and they carry total quark spin $S=\nicefrac{{1}}{{2}}$ and orbital angular momentum $L=1$ which are coupled to $J=\nicefrac{{1}}{{2}}$ and $\nicefrac{{3}}{{2}}$. The next cluster around 1.65 GeV contains a $N^{*}$ triplet, the $N(1650)\nicefrac{{1}}{{2}}^{-}$, $N(1700)\nicefrac{{3}}{{2}}^{-}$, and $N(1675)\nicefrac{{5}}{{2}}^{-}$ forming the $N(\mathbf{70},1^{-},\nicefrac{{3}}{{2}})$, and a $\Delta^{*}$ pair, the $\Delta(1620)\nicefrac{{1}}{{2}}^{-}$ and $\Delta(1700)\nicefrac{{3}}{{2}}^{-}$ forming the $\Delta(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$. Note that the first shell does not permit a $\Delta$ triplet.
The c.o.g. masses of these states are collected in Fig. 3.5. The small mass splittings within the two spin doublets and the spin triplet imply that the $\bm{L}\cdot\bm{S}$ coupling is small. The mass-square difference between the $N(\mathbf{70},1^{-},\nicefrac{{1}}{{2}})$ and the mean value of the five remaining states is 0.56(7) GeV2. In principle, the states with identical quantum numbers can mix, but the mixing angle is small: Experimentally and in calculations, the mixing angle is determined to be $31,7^{\circ}$ for the two $\nicefrac{{1}}{{2}}^{-}$ and +10 for the $\nicefrac{{3}}{{2}}^{-}$ states Isgur:1977ef . Here, we caution the reader: the mixing angle was derived assuming a small branching ratios of about 1% for $N(1650)\nicefrac{{1}}{{2}}^{-}\to N\eta$ from unreliable data on $\pi^{-}p\to\eta p$; the later determination was reported in Ref. Hunt:2018wqz . Recent determinations using precise data on $\eta$ photoproduction find, however, a branching ratio of 28(11)% A2:2017gwp and 33(4)% CBELSATAPS:2019ylw .
@@
Figure 3.5: Squared masses of stretched $N^{*}$’s (a) and $\Delta^{*}$’s (b) as a function of their spin $J$. c: C.o.g. masses (in MeV) for the baryon excitations in Fig. 3.2 up to the second excitation shell. d: Mass splittings in the second and third excitation shell expected in the quark model, see Refs. Isgur:1977ef , Bowler:1980pfa , Bowler:1981xh .
@@ 3.3.2   Second excitation shell
$(\mathbf{56},0^{+})$: The so-called radial excitations of the nucleon and $\Delta(1232)$ belong to the 56-plet, so they have symmetric orbital and spin-flavor wave functions and carry orbital angular momentum $L=0$. The low mass of 1366(3) MeV for the Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$ compared to the $N(1535)\nicefrac{{1}}{{2}}^{-}$ has led to speculations concerning its nature Burkert:2017djo . Its squared mass is 0.98(1) GeV2 above $M^{2}_{N}$. The $\Delta(1600)\nicefrac{{3}}{{2}}^{+}$ with a pole mass of 1550(15) MeV is the analogue of the Roper and was nicknamed ‘Doper’ by Nefkens (who also introduced the names ‘Loper’ for $\Lambda(1600)\nicefrac{{1}}{{2}}^{+}$, ‘Soper’ for $\Sigma(1660)\nicefrac{{1}}{{2}}^{+}$ and ‘Xoper’ for the not yet identified spin-$\nicefrac{{1}}{{2}}^{+}$ $\Xi$ resonance in his talk at MENU2004 Nefkens:2005dh ). The $\Delta(1600)\nicefrac{{3}}{{2}}^{+}-N(1440)\nicefrac{{1}}{{2}}^{+}$ mass-square difference of 0.54(4) GeV2 is compatible with the $\Delta(1232)-N(940)$ mass-square difference of 0.58 GeV2.
@@
$(\mathbf{70},0^{+})$: The remaining $\text{N}=2$ mass region is rather densely packed with baryon excitations. The $N(1710)\nicefrac{{1}}{{2}}^{+}$ has a pole mass of 1696(10) MeV. Given that its squared mass splitting with the Roper resonance of 1.01(4) GeV2 is of comparable size as that of the Roper versus the nucleon, it could be interpreted as the second radial excitation of the nucleon in a 56-plet. In quark models, however, the $N(1710)\nicefrac{{1}}{{2}}^{+}$ is found in a 70-plet, which has a spatial wave functions with mixed symmetry. In that case, both $\rho$ and $\lambda$ oscillators carry one angular momentum excitation and we expect a sizable fraction of all decays to proceed via an intermediate excitation, i.e., via $N(1440)\nicefrac{{1}}{{2}}^{+}$ or $N\sigma$. In the RPP these two decay modes are not listed as seen, but in a recent analysis the former branching ratio is reported with a branching ratio 22(6)% and the latter with 11(4)% CLAS:2024iir , Sarantsev:2025lik . Thus, in agreement with the quark model interpretation we consider the $N(1710)\nicefrac{{1}}{{2}}^{+}$ as the $N(\mathbf{70},0^{+},\nicefrac{{1}}{{2}})$ state. We then further expect a nucleon spin partner $N(\mathbf{70},0^{+},\nicefrac{{3}}{{2}})$ with $J=\nicefrac{{3}}{{2}}^{+}$ in this multiplet. This could be the $N^{\prime}(1720)\nicefrac{{3}}{{2}}^{+}$ suggested in a common photo- and electroproduction analysis of $p\pi^{+}\pi^{-}$ Mokeev:2015lda , which is not included in the RPP 2024. The remaining decuplet state $\Delta(\mathbf{70},0^{+},\nicefrac{{1}}{{2}})$ could be the 1* resonance $\Delta(1750)\nicefrac{{1}}{{2}}^{+}$ with a mass of 1770(30) MeV.
@@
$(\mathbf{56},2^{+})$: In the same mass region there are several further states: a spin doublet $N(\mathbf{56},2^{+},\nicefrac{{1}}{{2}})$ containing the $N(1720)\nicefrac{{3}}{{2}}^{+}$ and $N(1680)\nicefrac{{5}}{{2}}^{+}$ with c.o.g. mass 1689(13) MeV, and a spin quartet $\Delta(\mathbf{56},2^{+},\nicefrac{{3}}{{2}})$ with the $\Delta(1910)\nicefrac{{1}}{{2}}^{+}$, $\Delta(1920)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1905)\nicefrac{{5}}{{2}}^{+}$ and $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$ and c.o.g. mass 1859(20) MeV. We note that the $\Delta(1905)\nicefrac{{5}}{{2}}^{+}$ has an unexpectedly low pole mass of 1795(25) MeV. The mass square from $\Delta(\mathbf{56},2^{+},\nicefrac{{3}}{{2}})$ to $N(\mathbf{56},2^{+},\nicefrac{{1}}{{2}})$ is 0.60(8) GeV2 and $2\times$ 1.00(8) GeV2 to the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$.
@@ 3.3.3   The third shell
$(\mathbf{56},1^{-})$: There are five negative-parity states with masses that are difficult to reconcile in quark models. In Capstick:1986ter , Loring:2001kx , the lowest-mass $N^{*}$ doublet in the third shell is expected at about 1950 MeV and a $\Delta^{*}$ triplet at around 2100 MeV. Experimentally, we find a cluster of $N^{*}$’s forming a spin-doublet, the $N(1895)\nicefrac{{1}}{{2}}^{-}$ and $N(1875)\nicefrac{{3}}{{2}}^{-}$ with c.o.g. mass 1868(23) MeV, and the spin triplet $\Delta(1900)\nicefrac{{1}}{{2}}^{-}$, $\Delta(1940)\nicefrac{{3}}{{2}}^{-}$, and $\Delta(1930)\nicefrac{{5}}{{2}}^{-}$ with c.o.g. mass 1858(63) MeV. This would fit with the $N(\mathbf{56},1^{-},\nicefrac{{1}}{{2}})$ and $\Delta(\mathbf{56},1^{-},\nicefrac{{3}}{{2}})$ assignment. To explain its low mass, one of these states, the $N(1895)\nicefrac{{1}}{{2}}^{-}$, has been interpreted as a dynamically generated state with a two-pole structure Khemchandani:2013nma , Khemchandani:2020exc .
@@ 3.3.4   Beyond the third shell
In the negative-parity sector, the $\Delta(2350)\nicefrac{{5}}{{2}}^{-}$ and $\Delta(2400)\nicefrac{{9}}{{2}}^{-}$ could be naturally assigned to the $\Delta(\mathbf{56},3^{-},\nicefrac{{3}}{{2}})$ multiplet in the third excitation shell (see Fig. 3.2). However, the former 1* resonance is observed at 2400(125) MeV Cutkosky:1980rh , while for the latter 2* resonance two measurements have been reported, 2260(60) MeV Cutkosky:1980rh and 2458(2) MeV Ronchen:2022hqk . The unweighted mean of these three values is 2375(100) MeV, which is considerably higher than the masses of the other states in the third shell. Moreover, the masses of these two resonances fall onto a Regge trajectory with an offset compared to main trajectory given by the $\Delta(1600)\nicefrac{{3}}{{2}}^{+}-\Delta(1232)\nicefrac{{3}}{{2}}^{+}$ squared mass separation, as shown in the right panel of Fig. 3.6. We thus prefer an interpretation of these states as members of $\Delta(\mathbf{56},3^{-},\nicefrac{{3}}{{2}})$ in the fifth excitation shell ($\text{N}=5$). For the same reasons one can assign the 3* resonance $N(2600)\nicefrac{{11}}{{2}}^{-}$ to the multiplet $N(\mathbf{70},5^{-},\nicefrac{{1}}{{2}})$ in the fifth shell and the $\Delta(2750)\nicefrac{{13}}{{2}}^{-}$ with mass 2794(80) MeV to $\Delta(\mathbf{56},5^{-},\nicefrac{{3}}{{2}})$ in the seventh shell.
Concerning the positive-parity resonances, the $\Delta(2420)\nicefrac{{11}}{{2}}^{+}$, $\Delta(2300)\nicefrac{{9}}{{2}}^{+}$ and $\Delta(2390)\nicefrac{{7}}{{2}}^{+}$ have a c.o.g. mass of 2360(100) MeV and can be combined into a multiplet $\Delta(\mathbf{56},4^{+},\nicefrac{{3}}{{2}})$ in the fourth shell, where the $J=\nicefrac{{5}}{{2}}^{+}$ state is missing. In the nucleon sector, the RPP lists the 4* resonance $N(2220)\nicefrac{{9}}{{2}}^{+}$; the $N(2200)\nicefrac{{7}}{{2}}^{+}$ suggested by Hunt and Manley Hunt:2018wqz could serve as its spin partner to form a $N(\mathbf{56},4^{+},\nicefrac{{1}}{{2}})$ in the fourth shell. Finally, the $N(2700)\nicefrac{{13}}{{2}}^{+}$ and $\Delta(2950)\nicefrac{{15}}{{2}}^{+}$ are the highest-mass states included in the RPP; the Regge trajectories in Fig. 3.6 suggest to place the former in the sixth shell with $N(\mathbf{56},6^{+},\nicefrac{{1}}{{2}})$ and the latter in the eight shell with $\Delta(\mathbf{56},6^{+},\nicefrac{{3}}{{2}})$.
Table 3.2: Number of expected and (in parentheses) reported $N^{*}$ and $\Delta^{*}$ resonances with $J^{P}=\nicefrac{{1}}{{2}}^{\pm},\cdots,\nicefrac{{9}}{{2}}^{\pm}$ in the first three excitation shells. The $N^{\prime}(1720)\nicefrac{{3}}{{2}}^{+}$ suggested in electroproduction  Mokeev:2015lda and $N(1965)\nicefrac{{3}}{{2}}^{+}$ suggested in CLAS:2024iir , Sarantsev:2025lik are not included in the counting.
NN Shell $P$ $J=\frac{1}{2}$ $J=\frac{3}{2}$ $J=\frac{5}{2}$ $J=\frac{7}{2}$ $J=\frac{9}{2}$
1st - 2 (2) 2 (2) 1 (1)
@@
3rd - 3 (2) 5 (1) 4 (2) 2 (1) 1 (0)
Table 3.3: Difference in squared masses of radial excitations. In some cases the SU(6)U(6) multiplets are not identical on the left- and right-hand side; e.g., the $N(1440)\nicefrac{{1}}{{2}}^{+}$ is in a 56-plet but the $N(1710)\nicefrac{{1}}{{2}}^{+}$ in a 70-plet.
$X$ $Y$ $\delta M^{2}$ [GeV2]
NN $N(1440)\frac{1}{2}^{+}$ $0.98\pm 0.01$
$N(1440)\frac{1}{2}^{+}$ $N(1710)\frac{1}{2}^{+}$ $1±\pm0$.04
@@
[…]